OpenStudy (anonymous):

if you have frequency of 440 then what is the wavelength

OpenStudy (xishem):

The wave-speed equation relates the frequency and wavelength of an equation:\[s=\lambda\nu\]Where s is the propagation speed of the wave, lambda is the wavelength, and nu is the frequency.

OpenStudy (xishem):

In other words, it depends what the propagation speed of the wave is.

OpenStudy (anonymous):

We know that,velocity=wavelength*frequency. here wavelength=velocity/frequency=3*10000000000cm/sec/440cycles/second=0.068*1000000000cm/cycles

OpenStudy (xishem):

The problem doesn't specify that it's a wave of light — if it were, 440Hz is an incredibly low frequency for a light wave, and that's why I'm partially detracted from just assuming that it is light.

OpenStudy (unklerhaukus):

I think that in music the note A (above middle C) has frequency \(\nu=440 [\text{Hz}]\) the speed of sound in air is about \(s=340 [\text m/\text s]\) Use the wave-speed equation Xishem has provided, first rearrange it, to solve for \(\lambda\) (the wavelength) then just substitute in frequency and speed to get the units to simplify, use the fact that hertz are inverse seconds \([\text{Hz}]\equiv [\frac1{\text s}]\)

OpenStudy (anonymous):

A440 is the musical note A above middle C. It has a frequency of 440 Hz and is a general tuning standard for musical pitch. So you would use the speed of sound in air which is given by speed = 331.4 + 0.6*Temp = 343.6 m/s at 20 deg C Lamda = speed / frequency = 343.6 / 440 = 0.78 meters